Determine the equation of the line containing the points (0, 32) & (100, 212).... Write the equation in "y = mx + b" form..... also write equation in "Ax + By = C" form
Answer by Alan3354(31513)

slope m = diffy/diffx = 180/100 = 1.8
b = 32
--> y = 1.8x + 32
1.8x - y = -32
-------------
This the the conversion between Fahrenheit and Celsius, btw